# Geometric measures of CHSH quantum nonlocality: characterization, quantification, and comparison by distances and operations

### v3

@@ Abstract
We introduce a geometric framework for analyzing Bell nonlocality in Hilbert space, aiming to provide intrinsic measures of nonlocality based solely on the properties of quantum states for a given Bell inequality. Recognizing that nonlocality is inherently measurement-dependent, our approach focuses on fixing a specific Bell scenario—that is, a chosen number of measurements and outcomes per party, along with a specific Bell inequality—and, for each quantum state, maximizing the violation over all possible measurement choices within the given scenario. We then characterize the geometry of the corresponding set of local states, defined as those states for which no violation occurs for any choice of measurements. Within this framework, we define geometric and entropic measures of nonlocality based on the distance of a given quantum state from the set of local states, using quantities such as trace distance, Hilbert-Schmidt distance, and relative entropy. We first establish the general formalism, emphasizing the challenges posed by the NP-hardness of characterizing local sets for arbitrary Bell scenarios. We then specialize to the case of the Clauser-Horne-Shimony-Holt (CHSH) inequality in two-qubit systems, where the local set is fully characterized and an explicit geometric analysis is possible. In this setting, we derive geometric measures of nonlocality for Bell-diagonal and Werner states, showing that the closest local state to a Werner state remains a Werner state, and similarly for Bell-diagonal states. For each quantum states, our results offer a metric-based characterization of Bell nonlocality that is independent of specific experimental implementations and naturally extend to different Bell scenarios when the corresponding set of local states is known.
@@ VI.2.1 Hilbert-Schmidt distance
$\tilde{a}_{1}^{loc}=\frac{-a_{2}\pm\sqrt{a_{2}^{2}+4}}{2},$
$\tilde{a}_{2}^{loc}=\frac{-a_{1}\pm\sqrt{a_{1}^{2}+4}}{2},$
$\tilde{a}_{3}^{loc}=a_{3}.$ (70)
@@
$\mathcal{M}_{HS}(\rho_{BD})=\frac{1}{4}\left[(a_{1}-a_{2})^{2}+\frac{a_{1}^{2}+a_{2}^{2}}{2}+(a_{2}-a_{1})\bigg{(}\sqrt{a_{1}^{2}+4}-\sqrt{a_{2}^{2}+4}\bigg{)}+2\right].$ (71)
The remaining cases $|a_{2}^{loc}|,|a_{3}^{loc}|>|a_{1}^{loc}|$ and $|a_{1}^{loc}|,|a_{3}^{loc}|>|a_{2}^{loc}|$ can be immediately found by symmetry and the global minimum is determined over all possible cases. For the maximally nonlocal Bell states the minimum is attained when $|a_{1}^{loc}|=|a_{2}^{loc}|=|a_{3}^{loc}|\equiv w^{2}$, with $w$ the Werner parameter. The constraints in Eq. (69) thus reduce to $12w^{2}=$ 1, i.e. $w=1/\sqrt{2}$, and $\mathcal{M}_{HS}$ achieves its maximum $\mathcal{M}_{HS}^{max}$, Eq. (59). Using Eqs. (41) and (42) the transformation from the matrix elements $\{a_{i}\}$ to the eigenvalues $\{e_{i}\}$ of the density matrix follows immediately. The behavior of the normalized HS measure of CHSH nonlocality ~$\widetilde{\mathcal{M}}_{HS}=\mathcal{M}_{HS}/\mathcal{M}_{HS}^{max}$ for Bell-diagonal states as a function of the eigenvalues $e_{1}$, $e_{2}$ in the case of $e_{4}=0$ is reported in Fig. 3.

### v4

@@ Abstract
We introduce a geometric framework for the study of Bell nonlocality in Hilbert space, aiming to provide intrinsic measures based solely on the properties of quantum states for a given Bell inequality. Recognizing that nonlocality is inherently measurement-dependent, our approach focuses on fixing a specific Bell scenario—that is, a chosen number of measurements and outcomes per party, along with a specific Bell inequality—and, for each quantum state, maximizing the violation over all possible measurement choices within the given scenario. We then characterize the geometry of the corresponding set of local states, defined as those states for which no violation occurs for any choice of the measurements. Within this framework, we define geometric measures of nonlocality based on the distance of a given quantum state from the set of local states, using quantities such as the trace distance, the Hilbert-Schmidt distance, and the relative entropy. We first establish the general formalism, emphasizing the challenges posed by the NP-hardness of characterizing local sets for arbitrary Bell scenarios. We then specialize to the case of the Clauser-Horne-Shimony-Holt (CHSH) inequality in two-qubit systems, where the local set is fully characterized and an explicit geometric analysis is possible. In this setting, we derive geometric measures of nonlocality for Bell-diagonal and Werner states, showing that the local state closest to a Werner state is a Werner state, and analogously for Bell-diagonal states. For each quantum state, our results offer a metric-based characterization of Bell nonlocality that is independent of specific experimental implementations and naturally extend to different Bell scenarios whenever the corresponding set of local states is accessible.
@@ VI.2.1 Hilbert-Schmidt distance
$a_{1}^{loc}=\pm\frac{a_{1}}{\sqrt{a_{1}^{2}+a_{2}^{2}}}{2},$
$a_{2}^{loc}=\pm\frac{a_{2}}{\sqrt{a_{2}^{2}+a_{3}^{2}}},$
$a_{3}^{loc}=a_{3}.$ (70)
@@
$\mathcal{M}_{HS}(\rho_{BD})=\frac{3}{4}\sum\limits_{i=x,y,z}(a_{i}-a_{i}^{loc})^{2}=\frac{3}{4}\left(\sqrt{a_{1}^{2}+a_{2}^{2}}\mp 1\right)^{2}$ (71)
whose minimum clearly correspond to the minus sign (the plus sign in (71)), since $t_{x}^{2}+t_{y}^{2}\geq 1$. The remaining cases $|a_{2}^{loc}|,|a_{3}^{loc}|>|a_{1}^{loc}|$ and $|a_{1}^{loc}|,|a_{3}^{loc}|>|a_{2}^{loc}|$ can be immediately found by symmetry and the global minimum is determined over all possible cases. For the maximally nonlocal Bell states the minimum is attained when $|a_{1}^{loc}|=|a_{2}^{loc}|=|a_{3}^{loc}|\equiv w^{2}$, with $w$ the Werner parameter. The constraints in Eq. (69) thus reduce to $12w^{2}=$ 1, i.e. $w=1/\sqrt{2}$, and $\mathcal{M}_{HS}$ achieves its maximum $\mathcal{M}_{HS}^{max}$, Eq. (59). Note that this analytical expression is not always valid, as no constraint has been imposed on the positivity of the density operator associated with the set $\{a_{i}^{loc}\}$. Therefore, in order to use expression (71), one must also take such constraints into account. If the local state does not satisfy these constraints, the distance must be computed numerically. Using Eqs. (41) and (42) the transformation from the matrix elements $\{a_{i}\}$ to the eigenvalues $\{e_{i}\}$ of the density matrix follows immediately. The behavior of the normalized HS measure of CHSH nonlocality ~$\widetilde{\mathcal{M}}_{HS}=\mathcal{M}_{HS}/\mathcal{M}_{HS}^{max}$ for Bell-diagonal states as a function of the eigenvalues $e_{1}$, $e_{2}$ in the case of $e_{4}=0$ is reported in Fig. 3.